OpenStudy (anonymous):

Medal for fastest correct answer --> 3+5

OpenStudy (anonymous):

16

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do i have to fourier?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

8 ! =)

OpenStudy (anonymous):

how do you know when your bored...

OpenStudy (anonymous):

No, \[8!=1 \times 2 \times 3...\times 8\]

OpenStudy (anonymous):

@fandes